Bramble Corporation is a small wholesaler of gourmet food products. Data regarding the store's operations follow:

  • Sales are budgeted at $260,000 for November, $240,000 for December, and $230,000 for January.
  • Collections are expected to be 55% in the month of sale and 45% in the month following the sale.
  • The cost of goods sold is 80% of sales.
  • The company would like to maintain ending merchandise inventories equal to 70% of the next month's cost of goods sold. Payment for merchandise is made in the month following the purchase.
  • Other monthly expenses to be paid in cash are $23,200.
  • Monthly depreciation is $14,200.
  • Ignore taxes.
Balance Sheet
October 31
Assets
Cash $ 22,000
Accounts receivable 72,000
Merchandise inventory 145,600
Property, plant and equipment, net of $574,000 accumulated depreciation 1,096,000
Total assets $ 1,335,600
Liabilities and Stockholders' Equity
Accounts payable $ 256,000
Common stock 822,000
Retained earnings 257,600
Total liabilities and stockholders' equity $ 1,335,600

The difference between cash receipts and cash disbursements for December would be:

(A) $29,000

(B) $21,750

(C) $19,400

(D) $72,600
Answer #1

Cash receipt for December = 260000*45%+240000*55% = 249000

November purchase = 260000*80%+240000*80%*70%-145600 = 196800

Cash payment for December = 196800+23200 = 220000

Difference = 249000-220000 = 29000

So answer is a) $29000